# name=Sara Hooker

### v1

@@ Abstract
At face value, this essay is about understanding a fairly esoteric governance tool called compute thresholds. However, in order to grapple with whether these thresholds will achieve anything, we must first understand how they came to be. This requires engaging with a decades-old debate at the heart of computer science progress, namely, is “bigger always better?” Hence, this essay may be of interest not only to policymakers and the wider public but also to computer scientists interested in understanding the role of compute in unlocking breakthroughs. Does a certain inflection point of compute result in changes to the risk profile of a model? This discussion is increasingly urgent given the wide adoption of governance approaches that suggest greater compute equates with higher propensity for harm. Several leading frontier AI companies have released responsible scaling policies. Both the White House Executive Orders on AI Safety (EO) and the EU AI Act encode the use of FLOP or “floating-point operations” as a way to identify more powerful systems. What is striking about the choice of compute thresholds to-date is that no models currently deployed in the wild fulfill the current criteria set by the EO. This implies that the emphasis is often not on auditing the risks and harms incurred by currently deployed models – but rather is based upon the belief that future levels of compute will introduce unforeseen new risks. A key conclusion of this essay is that compute thresholds as currently implemented are shortsighted and likely to fail to mitigate risk. Governance that is overly reliant on compute fails to understand that the relationship between compute and risk is highly uncertain and rapidly changing. It also overestimates our ability to predict what abilities emerge at different scales. This essay ends with recommendations for a better way forward.
@@ 1 Understanding Risk
(a) The plague. An example where lack of medical knowledge precluded an understanding of what levers amplified risk.
(b) The great fire of London. An example where the sources of risk were known, but the proportional response was inadequate.
@@
Risk is a particularly challenging concept to formulate an effective governance response to, because it requires both 1) a successful estimate of the level and origins of risk to society and 2) aligning on a proportionate response. History is replete with examples where one or both requirements fail. For example, the large human toll incurred by the black death is a good example of the difficulty of estimating what vectors amplify risk, where inadequate medical knowledge in the 1300s led to a failure to identify rats as one of the main carriers of the disease (Benedictow, 2004). In other cases, the risk is well known yet the response is inadequate. In 1966, the famous London fire swept through the city and devastated over half of all buildings. This risk was well known by authorities, as London had experienced several major fires before 1666. However, hesitation from authorities to act quickly to contain the blaze doomed the city (Peter, 2002).
@@
With more powerful tools comes more possibility for misuse. This includes known harms including propensity for hallucinations (HAI, 2023; Economist, 2023; Kossen et al., 2024), disinformation and misinformation (Zhou et al., 2023; Zellers et al., 2019; Goldstein et al., 2023; Musser, 2023; Buchanan et al., 2021), bias (Wiggers, 2023; Hao, 2024) and toxicity (Pozzobon et al., 2023a; Üstün et al., 2024; Gehman et al., 2020). However, it also includes unknown risks incurred by further developing this technology, with researchers concerned by national security risks like biorisk (AISI, 2024; Mouton et al., 2024; OpenAI, 2024), cybersecurity threats (NCSC, 2024; Barrett et al., 2024; Fang et al., 2024a; Lohn & Musser, 2022) and loss of control (UK Government, 2021). Partly, the difficulty we face is how to balance this portfolio of risks and how to allocate limited resources between mitigation of both present and future possible harms.
@@
(a) Open Leaderboard Scores For Small Models (<13B) Over Time
(b) Open Leaderboard Scores for large models (>13B) that perform worse Small Models (<13B) Over Time.
@@
These are not isolated examples, but rather indicative of an overall trend where there is no guarantee larger models consistently outperform smaller models. Figure 3(b) plots the scores of models submitted to the Open LLM Leaderboard over the last two years. Here, we plot large models with more than 13 billion parameters whose leaderboard score is less than the top performing small model with less than 13 billion parameters. We observe that over time, more and more large models have been submitted that are outperformed by the best small model daily submission. To understand why this is the case, we must understand what key variables have been driving gains in performance over the last decade. In an era where there are diminishing returns for the amount of compute available (Lohn & Musser, 2022; Thompson et al., 2020), optimization and architecture breakthroughs define the rate of return for a given unit of compute. It is this rate of return which is most critical to the pace of progress and to the level of risk incurred by additional compute.
@@ 2.1 A shift in the relationship between compute and performance
Data quality reduces reliance on compute. Models trained on better data do not require as much compute. A large body of work has emerged which shows that efforts to better curate training corpus, including de-duping (Taylor et al., 2022; Kocetkov et al., 2022), data pruning (Marion et al., 2023; Singh et al., 2024a; Sorscher et al., 2023; Albalak et al., 2024; Tirumala et al., 2023; Chimoto et al., 2024) or data prioritization (Boubdir et al., 2023; Thakkar et al., 2023) can compensate for more weights. This suggests that the number of learnable parameters is not definitively the constraint on improving performance; investments in better data quality mitigate the need for more weights (Singh et al., 2024a; Penedo et al., 2023; Raffel et al., 2020; Lee et al., 2022). If the size of a training dataset can be reduced without impacting performance (Marion et al., 2023), training time is reduced. This directly impacts the number of training FLOP and means less compute is needed.
Figure 4: Bytes Magazine Cover, Volume 5, 1980. Compute is rarely the only determinant of progress. Data quality (Marion et al., 2023), instruction-finetuning (Singh et al., 2024a), preference training (Dang et al., 2024), retrieval augmented networks (Pozzobon et al., 2023b), enabled tool use (Qin et al., 2023), chain-of-thought reasoning (Wei et al., 2023), increased context-length (Xiong et al., 2023) are all examples of optimization techniques which add little or no training FLOP but result in significant gains in performance.
@@
Optimization breakthroughs compensate for compute. Progress over the last few years has been as much due to optimization improvements as it has been due to compute. This includes extending pre-training with instruction finetuning to teach models instruction following (Singh et al., 2024b), model distillation using synthetic data from larger more performant "teachers" to train highly capable, smaller "students" (Team et al., 2024b; Aryabumi et al., 2024), chain-of-thought reasoning (Wei et al., 2023; Hsieh et al., 2023), increased context-length (Xiong et al., 2023), enabled tool-use (Qin et al., 2023; Wang et al., 2023a), retrieval augmented generation (Pozzobon et al., 2023b; Lewis et al., 2020), and preference training to align models with human feedback (Dang et al., 2024; Ahmadian et al., 2024; Ouyang et al., 2022; Bai et al., 2022; Lee et al., 2023; Tunstall et al., 2023; Khalifa et al., 2021; Rafailov et al., 2023; Azar et al., 2023). All these techniques compensate for the need for weights or expensive prolonged training (Ho et al., 2024b). All things equal, these have been shown to dramatically improve model performance relative to a model trained without these optimization tricks given the same level of compute (Davidson et al., 2023; Hernandez & Brown, 2020; Erdil & Besiroglu, 2023; METR Team, ; Liu et al., 2024). In Figure 3(a), we plot the best daily 13B or smaller model submitted to the Open LLM Leaderboard over time. In a mere span of 2 years, the best-performing daily scores from small model went from an average of 38.59% across to an average of 77.15% across 2024 submissions. The takeaway is clear – smaller models with the same amount of capacity are becoming more and more performant.
@@
Architecture plays a significant role in determining scalability The introduction of a new architecture design can fundamentally change the relationship between compute and performance (Tay et al., 2022; Sevilla et al., 2022a; Ho et al., 2024a) and render any compute threshold that is set irrelevant. For example, the key breakthroughs in AI adoption around the world were the introduction of architectures like convolutional neural networks (CNNs) for vision (Ciresan et al., 2011; Krizhevsky et al., 2012; Szegedy et al., 2014) and Transformers for language modeling (Vaswani et al., 2023).
@@
While deep neural networks represent a huge step forward in performance for a given level of compute, what is often missed is that our architectures also represent the ceiling in what is achievable through scaling. While progress has revolved around deep neural networks for the last decade, there is much to suggest that the next significant gain in efficiency will require an entirely different architecture. Deep neural networks remain very inefficient as an algorithm. Our typical training regimes require that all examples are shown the same number of times during the training (Xue et al., 2023). All modern networks are trained based upon minimization of average error (Goodfellow et al., 2016). This means that learning rare artifacts requires far more training time or capacity due to the diluted signal of infrequent attributes relative to the most frequent patterns in the dataset (Achille et al., 2017; Jiang et al., 2020; Mangalam & Prabhu, 2019; Faghri et al., 2020; Frankle et al., 2020; Arpit et al., 2017). Small models are already good at learning the most frequent features, and most easy features and common patterns are learned early on training with much harder rare features learned in later stages (Agarwal & Hooker, 2020; Paul et al., 2021; Mangalam & Prabhu, 2019; Siddiqui et al., 2022; Abbe et al., 2021). When we radically scale the size of a model, we show the most gains in performance on are rare and underrepresented attributes in the dataset – the long-tail (Hooker et al., 2019, 2020). Put differently, scaling is being used to inefficiently learn a very small fraction of the overall training dataset. Our reliance on global updates also results in catastrophic forgetting, where performance deteriorates on the original task because the new information interferes with previously learned behavior (Mcclelland et al., 1995; Pozzobon et al., 2023b). All this suggests that our current architecture choices are probably not final and key disruptions lie ahead. This is likely to radically change any scaling relationships, in the same way it has done in the last decade. For example, it is unlikely any prediction of how compute scales based upon architectures before deep neural networks holds true post-2012 after the introduction of convolutional neural networks.
@@ 3.1 Challenges of using FLOP as a metric
Training FLOP doesn’t account for post-training leaps in performance Applying scrutiny and regulation based upon training FLOP ignores that a lot of compute can be spent outside of training to improve performance of a model. This can be grouped under “inference-time compute” and can result in large performance gains that dramatically increase the risk profile of a model. The limited work to-date which has evaluated a subset of ‘inference-time compute” improvements estimates these can impart gains between 5x and 20x of base level post-training performance (Davidson et al., 2023).“Inference-time compute” includes best-of-n sampling techniques (Team et al., 2024a), chain-of-thought reasoning (Wei et al., 2023; Hsieh et al., 2023; Wang et al., 2023c) and model distillation using synthetic data (Aryabumi et al., 2024; Shimabucoro et al., 2024; Üstün et al., 2024; Team et al., 2024a). All these techniques require more compute at test-time because of the need to perform more forward passes of the model to generate additional samples. However, these are not reflected in training time costs and indeed can often reduce the compute needed during training. For example, smaller, more performant models are often trained on smaller amounts of synthetic data from a highly performant teacher (Villalobos & Atkinson, 2023; Huang et al., 2022). These improvements dramatically improve performance but are currently completely ignored by compute thresholds since they don’t contribute to training FLOP.
@@
Figure 5: Bytes Magazine Cover, Volume 10, 1985. A key difficulty setting compute thresholds is that different domains and downstream tasks (language, vision, biology) demand very different levels of training compute, and so one compute threshold is not suitable to rule them all. This imposes more technical overhead on governments who must correctly set a hard-coded benchmark for each area. Only one domain specific compute threshold has been set to-date, by the EO for biological models. However, it has already been surpassed by several models who do not present very different profiles of risk from previous generations so may have been set too low. Setting different compute thresholds across domains can also invite gamification. For example, one loophole to avoid the lower EO compute threshold for biological models is to preserve biology specific training data at less than 50%. This would avoid classification as a biological model.
@@
Difficulty Tracking FLOP across model lifecycle. Increasingly, training a model falls into distinct stages that all confer different properties. For example, unsupervised pre-training dominates compute costs because the volume of data is typically in the trillions of tokens (Cottier, 2023; Heim, 2023). Following this, there is instruction finetuning, which confers the model the ability to follow instructions (Singh et al., 2024a) and then preference training (Aakanksha et al., 2024; Ahmadian et al., 2024; Bai et al., 2022; Ouyang et al., 2022; Lee et al., 2023; Tunstall et al., 2023; Khalifa et al., 2021; Rafailov et al., 2023; Azar et al., 2023), which aligns model performance with human values. Between each of these steps models are often released publicly (Üstün et al., 2024; Touvron et al., 2023; Aryabumi et al., 2024), meaning that developers can take a model from a different developer and continue optimizing. The models with the most downloads on platforms like HuggingFace are base models which are most conducive for continued pre-training. As sharing of models at different stages of the life-cycle becomes more common, so will difficulties in tallying FLOP across the entire model life-cycle. Furthermore, it may simply be infeasible to trace federated, decentralized training of models where hardware often belongs to many different participants and training is conducted in a privacy-preserving manner (Don-Yehiya et al., 2023; Borzunov et al., 2023; Yuan et al., 2023; Qin et al., 2024).
How to handle Mixture of Experts (MoEs) and classic ensembling? MoEs (Zadouri et al., 2023; Shazeer et al., 2018; Riquelme et al., 2021; Du et al., 2022; Fedus et al., 2022; Tan et al., 2024) are examples of adaptive compute – where examples are routed to different parts of a model. This type of architecture can often provide powerful efficiency gains, as despite a much larger overall architecture, only a subset of weights are activated for a given example. Current policy frameworks do clearly not specify how to handle Mixture of Experts (MoEs), which constitute some of the most highly performant systems currently deployed, such as Mixtral (Jiang et al., 2024) and the Gemini family of models (Team et al., 2024a). However, this raises important questions – should the compute for each expert be counted towards total FLOP, or only the FLOP used to train the subset of experts that are active at inference time? Given final performance depends on all experts in an MoE, a recommendation should be to include all FLOP in the final consideration, but this is currently under-specified. It also raises the question of how to treat new hybrid techniques which train several specialized experts and then both average parameters and utilize routing (Sukhbaatar et al., 2024).

### v2

@@ Abstract
At face value, this essay is about understanding a fairly esoteric governance tool called compute thresholds. However, in order to grapple with whether these thresholds will achieve anything, we must first understand how they came to be. To do so, we need to engage with a decades-old debate at the heart of computer science progress, namely, is bigger always better? Does a certain inflection point of compute result in changes to the risk profile of a model? Hence, this essay may be of interest not only to policymakers and the wider public but also to computer scientists interested in understanding the role of compute in unlocking breakthroughs. This discussion is timely given the wide adoption of compute thresholds in both the White House Executive Orders on AI Safety (EO) and the EU AI Act to identify more risky systems. A key conclusion of this essay is that compute thresholds as currently implemented are shortsighted and likely to fail to mitigate risk. The relationship between compute and risk is highly uncertain and rapidly changing. Relying upon compute thresholds overestimates our ability to predict what abilities emerge at different scales. This essay ends with recommendations for a better way forward.
@@ 1 Understanding Risk
(a) The plague. A lack of medical knowledge precluded an understanding of what levers amplified risk.
(b) The great fire of London. Here, the sources of risk were known yet the proportional response was inadequate.
@@
Risk is a particularly challenging concept to formulate an effective governance response to, because it requires both 1) a successful estimate of the level and origins of risk to society and 2) aligning on a proportionate response. History is replete with examples where one or both requirements fail. For example, the large human toll incurred by the black death is a good example of the difficulty of estimating what vectors amplify risk, where inadequate medical knowledge in the 1300s led to a failure to identify rats as one of the main carriers of the disease (Benedictow, 2004). In other cases, the risk is well known yet the response is inadequate. In 1666, the famous London fire swept through the city and devastated over half of all buildings. This risk was well known by authorities, as London had experienced several major fires before 1666. However, hesitation from authorities to act quickly to contain the blaze doomed the city (Peter, 2002).
@@
With more powerful tools comes more possibility for misuse. This includes known harms including hallucinations (HAI, 2023; Economist, 2023; Kossen et al., 2024), disinformation and misinformation (Zhou et al., 2023; Zellers et al., 2019; Goldstein et al., 2023; Musser, 2023; Buchanan et al., 2021), bias (Wiggers, 2023; Hao, 2024) and toxicity (Pozzobon et al., 2023a; Üstün et al., 2024; Gehman et al., 2020). However, it also includes unknown risks incurred by further developing this technology, with researchers concerned by national security risks like biorisk (AISI, 2024; Mouton et al., 2024; OpenAI, 2024), cybersecurity threats (NCSC, 2024; Barrett et al., 2024; Fang et al., 2024a; Lohn & Musser, 2022) and loss of control (UK Government, 2021). Partly, the difficulty we face is how to balance this portfolio of risks and how to allocate limited resources between mitigation of both present and future possible harms.
@@
(a) Open Leaderboard Scores For Small Models (<13B) Over Time
(b) Large models (>13B) that perform worse than Small Models (<13B)
@@
These are not isolated examples, but rather indicative of an overall trend where there is no guarantee larger models consistently outperform smaller models. Figure 3(b) plots the scores of models submitted to the Open LLM Leaderboard (Beeching et al., 2023) over the last two years. Here, we plot large models with more than 13 billion parameters whose leaderboard score is less than the top performing small model with less than 13 billion parameters. We observe that over time, more and more large models have been submitted that are outperformed by the best small model daily submission.
To understand why this is the case, we must understand what key variables have been driving gains in performance over the last decade. In an era where there are diminishing returns for the amount of compute available (Lohn & Musser, 2022; Thompson et al., 2020), optimization and architecture breakthroughs define the rate of return for a given unit of compute. It is this rate of return which is most critical to the pace of progress and to the level of risk incurred by additional compute.
@@ 2.1 A shift in the relationship between compute and performance
Models trained on better data do not require as much compute. A large body of work has emerged which shows that efforts to better curate training corpus, including de-duping (Taylor et al., 2022; Kocetkov et al., 2022), data pruning (Marion et al., 2023; Singh et al., 2024a; Sorscher et al., 2023; Albalak et al., 2024; Tirumala et al., 2023; Chimoto et al., 2024) or data prioritization (Boubdir et al., 2023; Thakkar et al., 2023) can compensate for more weights. This suggests that the number of learnable parameters is not definitively the constraint on improving performance; investments in better data quality mitigate the need for more weights (Singh et al., 2024a; Penedo et al., 2023; Raffel et al., 2020; Lee et al., 2022). If the size of a training dataset can be reduced without impacting performance (Marion et al., 2023), training time is reduced. This directly impacts the number of training FLOP and means less compute is needed.
2.1.2 Optimization breakthroughs compensate for compute.
@@
Progress over the last few years has been as much due to optimization improvements as it has been due to compute. This includes extending pre-training with instruction finetuning to teach models instruction following (Singh et al., 2024b), model distillation using synthetic data from larger more performant "teachers" to train highly capable, smaller "students" (Team et al., 2024b; Aryabumi et al., 2024), chain-of-thought reasoning (Wei et al., 2023; Hsieh et al., 2023), increased context-length (Xiong et al., 2023), enabled tool-use (Qin et al., 2023; Wang et al., 2023a), retrieval augmented generation (Pozzobon et al., 2023b; Lewis et al., 2020), and preference training to align models with human feedback (Dang et al., 2024; Ahmadian et al., 2024; Ouyang et al., 2022; Bai et al., 2022; Lee et al., 2023; Tunstall et al., 2023; Khalifa et al., 2021; Rafailov et al., 2023; Azar et al., 2023).
All these techniques compensate for the need for weights or expensive prolonged training (Ho et al., 2024b).All things equal, these have been shown to dramatically improve model performance relative to a model trained without these optimization tricks given the same level of compute (Davidson et al., 2023; Hernandez & Brown, 2020; Erdil & Besiroglu, 2023; METR Team, 2023; Liu et al., 2024). In Figure 3(a), we plot the best daily 13B or smaller model submitted to the Open LLM Leaderboard over time. In a mere span of 2 years, the best-performing daily scores from small model went from an average of 38.59% across to an average of 77.15% across 2024 submissions. The takeaway is clear – smaller models with the same amount of capacity are becoming more and more performant.
2.1.3 Architecture plays a significant role in determining scalability
@@
The introduction of a new architecture design can fundamentally change the relationship between compute and performance (Tay et al., 2022; Sevilla et al., 2022a; Ho et al., 2024a) and render any compute threshold that is set irrelevant. For example, the key breakthroughs in AI adoption around the world were the introduction of architectures like convolutional neural networks (CNNs) for vision (Ciresan et al., 2011; Krizhevsky et al., 2012; Szegedy et al., 2014) and Transformers for language modeling (Vaswani et al., 2023).
While deep neural networks represent a huge step forward in performance for a given level of compute, what is often missed is that our architectures also represent the ceiling in what is achievable through scaling.
Figure 4: Bytes Magazine Cover, Volume 5, 1980. Compute is rarely the only determinant of progress. Data quality, instruction-finetuning, preference training, retrieval augmented networks, enabled tool use, chain-of-thought reasoning, increased context-length are all examples of optimization techniques which add little or no training FLOP but result in significant gains in performance.
@@
While progress has revolved around deep neural networks for the last decade, there is much to suggest that the next significant gain in efficiency will require an entirely different architecture. Deep neural networks remain very inefficient as an algorithm. Our typical training regimes require that all examples are shown the same number of times during the training (Xue et al., 2023). All modern networks are trained based upon minimization of average error (Goodfellow et al., 2016). This means that learning rare artifacts requires far more training time or capacity due to the diluted signal of infrequent attributes relative to the most frequent patterns in the dataset (Achille et al., 2017; Jiang et al., 2020; Mangalam & Prabhu, 2019; Faghri et al., 2020; Frankle et al., 2020; Arpit et al., 2017). Small models are already good at learning the most frequent features, and most easy features and common patterns are learned early on training with much harder rare features learned in later stages (Agarwal & Hooker, 2020; Paul et al., 2021; Mangalam & Prabhu, 2019; Siddiqui et al., 2022; Abbe et al., 2021).
When we radically scale the size of a model, we show the most gains in performance are on rare and underrepresented attributes in the dataset – the long-tail (Hooker et al., 2019, 2020). Put differently, scaling is being used to inefficiently learn a very small fraction of the overall training dataset. Our reliance on global updates also results in catastrophic forgetting, where performance deteriorates on the original task because the new information interferes with previously learned behavior (Mcclelland et al., 1995; Pozzobon et al., 2023b). All this suggests that our current architecture choices are probably not final and key disruptions lie ahead. This is likely to radically change any scaling relationships, in the same way it has done in the last decade. For example, it is unlikely any prediction of how compute scales based upon architectures before deep neural networks holds true post-2012 after the introduction of convolutional neural networks.
@@ 3.1 Challenges of using FLOP as a metric
3.1.1 Training FLOP doesn’t account for post-training leaps in performance
Applying scrutiny and regulation based upon training FLOP ignores that a lot of compute can be spent outside of training to improve performance of a model. This can be grouped under “inference-time compute” and can result in large performance gains that dramatically increase the risk profile of a model. The limited work to-date which has evaluated a subset of ‘inference-time compute” improvements estimates these can impart gains between 5x and 20x of base level post-training performance (Davidson et al., 2023).“Inference-time compute” includes best-of-n sampling techniques (Team et al., 2024a), chain-of-thought reasoning (Wei et al., 2023; Hsieh et al., 2023; Wang et al., 2023c) and model distillation using synthetic data (Aryabumi et al., 2024; Shimabucoro et al., 2024; Üstün et al., 2024; Team et al., 2024a). All these techniques require more compute at test-time because of the need to perform more forward passes of the model to generate additional samples. However, these are not reflected in training time costs and indeed can often reduce the compute needed during training. For example, smaller, more performant models are often trained on smaller amounts of synthetic data from a highly performant teacher (Villalobos & Atkinson, 2023; Huang et al., 2022). These improvements dramatically improve performance but are currently completely ignored by compute thresholds since they don’t contribute to training FLOP.
@@ 3.1.1 Training FLOP doesn’t account for post-training leaps in performance
3.1.2 Difficulty Tracking FLOP across model lifecycle.
Increasingly, training a model falls into distinct stages that all confer different properties. For example, unsupervised pre-training dominates compute costs because the volume of data is typically in the trillions of tokens (Cottier, 2023; Heim, 2023). Following this, there is instruction finetuning, which confers the model the ability to follow instructions (Singh et al., 2024a) and then preference training (Aakanksha et al., 2024; Ahmadian et al., 2024; Bai et al., 2022; Ouyang et al., 2022; Lee et al., 2023; Tunstall et al., 2023; Khalifa et al., 2021; Rafailov et al., 2023; Azar et al., 2023), which aligns model performance with human values. Between each of these steps models are often released publicly (Üstün et al., 2024; Touvron et al., 2023; Aryabumi et al., 2024), meaning that developers can take a model from a different developer and continue optimizing. The models with the most downloads on platforms like HuggingFace are base models which are most conducive for continued pre-training. As sharing of models at different stages of the life-cycle becomes more common, so will difficulties in tallying FLOP across the entire model life-cycle. Furthermore, it may simply be infeasible to trace federated, decentralized training of models where hardware often belongs to many different participants and training is conducted in a privacy-preserving manner (Don-Yehiya et al., 2023; Borzunov et al., 2023; Yuan et al., 2023; Qin et al., 2024).
3.1.3 How to handle Mixture of Experts (MoEs) and classic ensembling?
MoEs (Zadouri et al., 2023; Shazeer et al., 2018; Riquelme et al., 2021; Du et al., 2022; Fedus et al., 2022; Tan et al., 2024) are examples of adaptive compute – where examples are routed to different parts of a model. This type of architecture can often provide powerful efficiency gains, as despite a much larger overall architecture, only a subset of weights are activated for a given example. Current policy frameworks do clearly not specify how to handle Mixture of Experts (MoEs), which constitute some of the most highly performant systems currently deployed, such as Mixtral (Jiang et al., 2024) and the Gemini family of models (Team et al., 2024a). However, this raises important questions – should the compute for each expert be counted towards total FLOP, or only the FLOP used to train the subset of experts that are active at inference time? Given final performance depends on all experts in an MoE, a recommendation should be to include all FLOP in the final consideration, but this is currently under-specified. It also raises the question of how to treat new hybrid techniques which train several specialized experts and then both average parameters and utilize routing (Sukhbaatar et al., 2024).